Calculate the force of gravity that the Earth (mass 6.0 x 1024 kg) and the Moon (mass 7.4 x 1022 kg) exert on each other. The average Earth-Moon distance is 3.8 x 108 m.

1
Expert's answer
2021-11-02T10:07:32-0400

Given the that:

Mass of Earth="6.0 \u00d710^{24}kg"

Mass of Moon="7.4 \u00d710^{22} kg"

The average Earth to Moon distance is = "3.8 \u00d7 10^8 m"

Gravitational constant G= "6.67\u00d710^{-11}Nm^2\/kg^2"

"F=\frac{Gm_1m_2}{d^2}"

"F= \frac {6.67×10^{-11}Nm^2/kg^2×6.0 ×10^{24}kg×7.4 ×10^{22} kg}{3.8 × 10^8 m}=20.1×10^{19} N"

Therefore the for between Earth and Moon ="20.1\u00d710^{19}N"